OpenStudy (loser66):

Derivative checking Find derivative \(\dfrac{x}{x+(c/x)}\) Need help on figure out the mistake

OpenStudy (loser66):

\(f(x)=\dfrac{x}{x+cx^{-1}}\\f'(x) = \dfrac{(x+c/x)-(x) (1-cx^{-2})}{(x+(c/x))^2}=\dfrac{2c}{x(x+(c/x))^2}\)

OpenStudy (solomonzelman):

perhaps it would be better to simplify. \(\large\color{black}{ \displaystyle \frac{ x }{x+\frac{c}{x}} }\) \(\large\color{black}{ \displaystyle \frac{ x }{\frac{x^2}{x}+\frac{c}{x}} }\) \(\large\color{black}{ \displaystyle \frac{ x }{\frac{x^2+c}{x}} }\) \(\large\color{black}{ \displaystyle \frac{ x^2 }{x^2+c} }\).

OpenStudy (loser66):

but other student gave me the solution above and I didn't find out what is wrong with it.

OpenStudy (loser66):

oh, I found it out.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{ (x+\frac{c}{x})-x(1-\frac{c}{x^2})}{(x+\frac{c}{x})^2} }\) \(\large\color{black}{ \displaystyle \frac{ x+\frac{c}{x}-x+\frac{c}{x}}{(x+\frac{c}{x})^2} }\) \(\large\color{black}{ \displaystyle \frac{ \frac{2c}{x}}{(x+\frac{c}{x})^2} }\) \(\large\color{black}{ \displaystyle \frac{2c}{x(x+\frac{c}{x})^2} }\)

OpenStudy (loser66):

Actually, I made mistake at somewhere.

OpenStudy (solomonzelman):

yes, you are lacking x on bottom

OpenStudy (loser66):

Nothing is wrong, the answers are same

OpenStudy (solomonzelman):

I would rather work out (x^2)/(x^2+c)

OpenStudy (loser66):

Thanks a lot

OpenStudy (solomonzelman):

just as a matter of preference

OpenStudy (solomonzelman):

np